A wave is modeled with the function y(x,t) = (0.8m) cos((1rad/m)x − (0.62rad/s)t + π/3). Find the wave speed.


1
Expert's answer
2022-05-22T14:38:07-0400

The wave equation is given by


"y = A\cos((2\pi/\lambda)x − \omega t + \phi_0)\\
y= (0.8m) \cos((1rad/m)x − (0.62rad/s)t + π/3)"

The wave speed

"v=f\lambda=\omega\lambda/2\pi\\
=0.62*1=0.62\:\rm m/s"